# Optimizing Retinal Prosthetic Stimuli with Conditional Invertible Neural Networks

### v1

@@ 2.3 Conditional Invertible Neural Network
We first discuss the maximum likelihood training for non-conditional INNs. With $\mathbf{x}$ and $\mathbf{z}$ sharing the identical dimensionality, the bijective normalizing flow $h$ can be trained by maximizing likelihood of $\mathbf{x}$, which is equivalent to minimizing the Negative Log-Likelihood (NLL) loss:
$\mathcal{L}_{\mathrm{NLL}}=-\log(p(\mathbf{x}))$ $=-\log(\mathcal{N}(h(\mathbf{x});\mathbf{0},\mathbf{I})))-\log\left|\det\mathbf{J}_{h}(\mathbf{x})\right|$
$\simeq\frac{1}{2}\|h(\mathbf{x})\|_{2}^{2}-\log\left|\det\mathbf{J}_{h}(\mathbf{x})\right|.$ (4)
Within the local coupling blocks [9], conditioning features $\mathbf{c}$ can be integrated into both directions of the network by concatenating to the subfunctions with $\mathbf{z}=h(\mathbf{x};\mathbf{c})$ and $\mathbf{x}=h^{-1}(\mathbf{z};\mathbf{c})$, forming Conditional Invertible Neural Networks (cINNs) [2]. Accordingly, the NLL loss for cINN maximum likelihood training is formulated as $\mathcal{L}_{\mathrm{NLL}}\simeq\frac{1}{2}\|h(\mathbf{x};\mathbf{c})\|_{2}^{2}-\log\left|\det\mathbf{J}_{h}(\mathbf{x})\right|$. As shown in Fig. 2b, we train the cINN by minimizing the NLL loss in an unsupervised manner and use the predicted percept during the training and the MNIST target during the evaluation as the conditioning. The same invertible neural network architecture is used as in the INN experiments.
@@ 3.1 Axon Map Model
We create a virtual square retinal implant with a $9\times 9$ electrode grid (Fig. 1) and utilize the physiologically validated Axon Map Model [5] implemented in pulse2percept [4]. The percept intensity at location $(x,y)$ is calculated by:
$I(x,y)=\exp\left(-\frac{(x-x_{\text{stim}})^{2}+(y-y_{\text{stim}})^{2}}{2\rho^{2}}-\frac{(x-x_{\text{soma}})^{2}+(y-y_{\text{soma}})^{2}}{2\lambda^{2}}\right),$ (5)
with exponential decays in terms of the distances to the stimulus at $(x_{\text{stim}},y_{\text{stim}})$ and to the elongated soma at $(x_{\text{soma}},y_{\text{soma}})$, controlled by the constants $\rho$ and $\lambda$, respectively [4, 5]. In this paper, we select realistic and challenging $\rho=400\,\mu m$ and $\lambda=1550\,\mu m$ for all experiments (cf. [5]).
@@ 3.2 Datasets
In order to achieve generalizability, models are trained using 10,000,000 random stimulus-percept pairs and then evaluated with the MNIST test set [18]. The stimulus always has a resolution $9\times 9$ pixels (i.e., the number of the electrodes) and we experiment with both $9\times 9$ and $28\times 28$ pixels for percepts and the MNIST targets. Each stimulus-percept pair of the training set is generated by uniformly sampling in $[-1,+1]$ for each electrode and the prediction of pulse2percept with Axon Map Model. Additionally, for each stimulus, only a random percentage of electrodes are activated while the rest are set to zero. Examples of stimulus-percept pairs are showcased in Fig. 5 in Appendix F.
@@ 3.3 Implementation Details and Preliminary Remarks
We compare five different approaches to optimize retinal prosthetic stimuli. The visualization of the network architectures are depicted in Fig. 8 in Appendix H.
@@
Neural network. For the non-invertible neural network models, we use a residual architecture with blocks of convolutional layers, batch normalizations [14], and ELU activations [7].
Invertible neural network. We use the open-source library FrEIA [1] for the implementation of the invertible network. Each INN layer is formed by a random permutation layer, a GLOW coupling block [1, 15], and an ActNorm layer [15]. The subfunction inside the coupling block contains two ReLU-activated linear layers with a hidden dimension of 512. A stack of 9 INN layers is empirically found to be sufficient and no additional performance gains are observed beyond that. During the training, we try to define either stimulus-to-percept or percept-to-stimulus as the forward process. We compare the MSE loss of both sides and find that defining $h$ (Eq. 2) as the percept-to-stimulus mapping makes the inverse function $h^{-1}$ easier to train, as shown in Fig. 7 in Appendix G. We thus report the results based on this setup in Table 1.
Conditional invertible neural network. The cINN architecture follows the design of the INN, with the additional conditioning to the subfunction at each layer [2]. The percept or the MNIST target is concatenated as the conditioning during the training or the inference, respectively (Fig. 2b). The conditional input is first transformed with a learnable linear layer to a size of $9\times 9$ before being passed to the subfunction. As introduced in Section 2.3, the cINN is trained in an unsupervised manner only with the NLL loss (Eq. 2.3).
(a) Correlation between likelihood and MSE for MNIST and random samples.
(b) Improved MSE with the optimized $\mathbf{z}$ for the 50 worst MNIST samples.
Figure 3: Investigation of log-likelihood and MSE loss in cINNs.
@@ 3.4 Likelihood and MSE
We illustrate whether the maximum likelihood training in the cINN approach leads to decrease the MSE loss [29]. The likelihood of the optimized stimulus depends on the likelihood of the latent $\mathbf{z}$ (Eq. 2.3). Fig. 2(a) shows the relationship between the log-likelihood of optimized stimulus (with $\mathbf{z}=\mathbf{0}$) and the MSE between the predicted percept and MNIST target (see Fig. 2b). For both random and MNIST samples, the correlation between the log-likelihood and the MSE can be observed. While setting $\mathbf{z}=\mathbf{0}$ should already give a good result, we further investigate if better $\mathbf{z}\neq\mathbf{0}$ can be found to yield higher likelihood and thus lower MSE. We select the worst 50 MNIST examples with the highest MSE and modify the latent $\mathbf{z}$ to maximize the total likelihood of optimized stimulus using gradient descent. Fig. 2(b) shows that the MSE is indeed decreased for most of the samples. Further optimization w.r.t. $\mathbf{z}$ can be addressed in future work.
Figure 4: Visualization of the reconstruction quality with five approaches. (I) denotes the original MNIST targets. (II) denotes the predicted percepts of $28\times 28$ pixels based on (III) optimized stimuli of $9\times 9$ electrodes.
@@ 3.5 Reconstruction Quality
The visualization of the reconstruction quality with a percept resolution of $28\times 28$ pixels is illustrated in Fig. 4. We argue that negative stimulations (in red) can counteract the surrounding or elongated phosphenes, which is not applicable in downsampling. The predicted percepts (II) from the cINN show clearer turnings (e.g., the bottom left part in the image of 2) and cleaner corners of the visual field. The advantages can be confirmed by quantitative assessments. Four evaluation metrics are reported in Table 1, namely Mean Absolute Error (MAE), Structural Similarity Index Measure (SSIM) [32], Peak Signal-to-Noise Ratio (PSNR), and classification accuracy with a pre-trained classifier (ACC) [12, 34]. With an increased percept resolution, cINN outperforms other approaches in various metrics. The number of the parameters of each model is listed in Table 1. The supervised models (linear, non-invertible and invertible neural networks) are susceptible to a higher resolution due to the fully-connected layers in the architectures, while this dense layer is only used to concatenate the target or percept to the subfunctions in the cINN approach.
@@ 4.1 Limitations
Our work relies on a pre-defined Axon Map Model with fixed $\rho$ and $\lambda$ for simulation. However, these parameters are largely patient- and implant-dependent [6] and not facile to determine in real-world scenarios (cf. [10, 11]). Additionally, we attempted to alter the dimensionality either on the input or output side and applied the MMD loss in INNs, which did not result in performance improvement. Finally, the pre-computed random samples, as illustrated in Fig. 2(a), are not sufficiently general, making the study of more complex datasets a challenge.
Table 1: Reconstruction quality evaluation from the predicted percepts to the original visual signals based on the optimized stimuli. Five approaches are compared with two resolutions (Res.) of the original MNIST input image of $9\times 9$ and $28\times 28$ pixels. Four evaluation metrics are reported, namely Mean Absolute Error (MAE), Structural Similarity Index Measure (SSIM), Peak Signal-to-Noise Ratio (PSNR), and classification accuracy with a pre-trained classifier (ACC). The number of the parameters of each model is listed. The top results are marked in red, while the second-best results are marked in blue.
 Methods    Res.
MAE $\downarrow$
SSIM $\uparrow$
PSNR $\uparrow$
ACC $\uparrow$
#Param $\downarrow$
down 0.1682 0.7248 30.4149 0.4272 1
linear 9 0.1657 0.7246 30.6877 0.5608 6,561
NN $\times$ 0.1387 0.7799 31.9620 0.5572 946,512
INN 9 0.1785 0.7088 30.5656 0.4892 1,866,270
cINN 0.1373 0.7651 30.8940 0.5480 1,936,710
down 0.3432 0.3784 25.0973 0.2884 1
linear 28 0.3279 0.4770 25.1453 0.5884 63,504
NN $\times$ 0.3340 0.4901 25.2400 0.6192 4,270,900
INN 28 0.3231 0.4937 25.2729 0.5764 18,063,390
cINN 0.3136 0.4943 25.5161 0.5932 2,449,197
@@ 4.2 Outlook
We expect our future studies to progress in two directions. First, supervised training of INN [3] can be further explored using different invertible architecture designs and training strategies. Second, simulation experiments can be conducted on more complex datasets using other computational models to optimize stimuli.
@@ 5.0.1 Acknowledgements
Appendix F Visualization
Figure 5: Examples of random samples from $9\times 9$ stimuli to both $9\times 9$ and $28\times 28$ percepts using Axon Map Model [5] implemented in pulse2percept [4].
Figure 6: Visualization of the reconstruction quality with five approaches. (I) denotes the original MNIST targets. (II) denotes the predicted percepts of $28\times 28$ pixels based on (III) optimized stimuli of $9\times 9$ electrodes. Continued from Fig. 4.
Appendix G MSE Loss in INN Trainings
Figure 7: MSE loss values of the INN training [3] for percept-to-stimulus being the forward process (left) and for stimulus-to-percept being the forward process (right). We observe that the MSE loss values of inverse processes are always larger than the supervised forward processes, which implies that the invertibility using a supervised INN is not guaranteed. We report the results of the left setup in Table 1, as the optimized stimulus is desired.
Appendix H Invertible Neural Network Architecture
Figure 8: We follow [2] to design the INN and cINN architectures with GLOW coupling layers [15]. The conditional input $\mathbf{c}$ in both forward and inverse GLOW coupling blocks is the only difference between the INN and the cINN approach. The affine coupling layer [9, 15] is implemented with GLOWCouplingBlock in [1].

### v2

@@ 3.3 Conditional Invertible Neural Network
Within the local coupling blocks [14], conditioning features $\mathbf{c}$ can be integrated into both directions of the network by concatenating them in the subfunctions with $\mathbf{z}=h(\mathbf{x};\mathbf{c})$ and $\mathbf{x}=h^{-1}(\mathbf{z};\mathbf{c})$, forming Conditional Invertible Neural Networks (cINNs) [4, 2]. This is similar to other conditional generative models, e.g., conditional variational auto-encoder (cVAE) [49] and conditional generative adversarial network (cGAN) [27]. Accordingly, the NLL loss for cINN maximum likelihood training is formulated as
$\mathcal{L}_{\mathrm{NLL}}\simeq\frac{1}{2}\|h(\mathbf{x};\mathbf{c})\|_{2}^{2}-\log\left|\det\mathbf{J}_{h}(\mathbf{x})\right|\,.$
As shown in Figure 3(b), we train the cINN by minimizing the NLL loss in an unsupervised manner and use the predicted percept during the training and the MNIST target during the evaluation as the conditioning. The same invertible neural network architecture is used as in the INN experiments. Unlike INN-MMD, which is constrained by a hybrid loss (combining supervised MSE and unsupervised MMD), cINN explicitly uses the conditioning as a reference to guide the stimulus with the target. A summary of notations for retinal implants from Section 3.1 is listed in Table 1, along with their corresponding contexts in INN-MMD and cINN.
(a) Forward process: percept-to-stimulus.
(b) Forward process: stimulus-to-percept.
Figure 5: MSE loss values of the INN-MMD training [3] (a) for percept-to-stimulus being the forward process and (b) for stimulus-to-percept being the forward process. We observe that the MSE loss values of inverse processes are always larger than the supervised forward processes, which implies that the invertibility using a supervised INN is not guaranteed. We report the results of the setup (a) in Table 2, as the optimized stimulus is desired.
@@ 4.1 Axon Map Model
In our in silico experiments, we create a virtual square retinal implant with a $9\times 9$ electrode grid (Figure 1(b)) and utilize the physiologically validated Axon Map Model [8], implemented in pulse2percept [7]. The phosphenes elicited by the axons, somas, or dendrites of the retinal ganglion cells have different shapes, as demonstrated in 1999 by Greenberg et al. [20]. The recent Axon Map Model [8] aligns with these findings by modelling the drawings from the subjects with retinal implants, which is one of the available models in pulse2percept [7] along with such as Scoreboard Model [8] and Biphasic Axon Map Model [17]. The predicted percept intensity at location $(x,y)$ is calculated by:
$\begin{split}I(x,y)=\exp&\left(-\frac{(x-x_{\text{stim}})^{2}+(y-y_{\text{stim}})^{2}}{2\rho^{2}}\right.\\ &\left.-\frac{(x-x_{\text{soma}})^{2}+(y-y_{\text{soma}})^{2}}{2\lambda^{2}}\right),\end{split}$ (18)
with exponential decays in terms of the distances to the stimulus at $(x_{\text{stim}},y_{\text{stim}})$ and to the elongated soma at $(x_{\text{soma}},y_{\text{soma}})$, controlled by the constants $\rho$ and $\lambda$, respectively (see Figure 1(b)). In this paper, we set the constants to the realistic and challenging values $\rho=400\,\mu m$ and $\lambda=1550\,\mu m$ for all experiments (cf. [8]).
@@ 4.2 Datasets
In order to achieve generalizability, models are trained using 10 million random stimulus-percept pairs and then evaluated with the MNIST test set [32]. We conduct all experiments using a batch size of 1024. The stimulus always has a resolution $9\times 9$ pixels (i.e., the number of the electrodes) and we experiment with both $9\times 9$ and $28\times 28$ pixels for percepts and the MNIST targets. Each stimulus-percept pair of the training set is generated by uniformly sampling in $[-3,+3]$ for each electrode. We select this intensity range to amplify the MSE values for particularly inaccurate predictions, when the difference between the predicted and ground-truth values exceeds 1. Negative voltage values in this range are beneficial in inhibiting phosphene elongation, which effectively designates the corresponding electrodes as return electrodes, as described in [54, 57]. We visualize stimuli and percepts with negative values in Figure 7, which do not result in any stimulated phosphenes though (similar to the neuron’s refractory period). Additionally, for each stimulus, only a random percentage of electrodes are activated while the rest are set to zero. Examples of stimulus-percept pairs are showcased in Figure 7(a).
(a) Correlation between log Jacobian determinant and MSE for MNIST and random samples.
(b) Improved MSE with the optimized $\mathbf{z}$ w.r.t. log Jacobian determinant for the 50 worst MNIST samples.
Figure 6: Investigation of likelihood (using log Jacobian determinant as in Equation 7) and MSE loss with cINNs. The y-axis is plotted on a logarithmic scale for better visualization.
@@ 4.3 Implementation Details and Preliminary Remarks
We compare five different approaches to optimize retinal prosthetic stimuli. The visualization of the invertible neural network architectures are depicted in Figure 4.
@@
Neural network. For the non-invertible neural network models, we use a feed-forward architecture consisting of three convolutional layers, each followed by batch normalization [26] and ELU activation [11]. The network is bookended by two linear layers at the input and output stages. Similar to the linear model, the feed-forward neural network is trained to map percepts to stimuli.
Invertible neural network with MMD. We use the open-source library FrEIA [1] for the implementation of the invertible network. Each INN layer is formed by a random permutation layer, Glow coupling block [1, 29], and an ActNorm layer [29]. The subfunction inside the coupling block contains two ReLU-activated linear layers with a hidden dimension of 512 and outputs both $s$ and $t$ simultaneously (Figure 4). A stack of 9 INN layers is empirically found to be sufficient and no additional performance gains are observed beyond that. The affine coupling layer [14, 29] is implemented using GLOWCouplingBlock in [1]. In contrast to AllInOneBlock in [1], GLOWCouplingBlock does not include ActNorm or invertible $1\times 1$ convolutions [29]. We also attempted to alter the dimensionality of the latent concatenations (Gaussian distribution in the forward path and zero paddings in the reverse path) either on the input or output side and applied the MMD loss in INNs, which did not result in performance improvement. During the training, we try to define either stimulus-to-percept or percept-to-stimulus as the forward process. We compare the MSE loss of both mappings and find that defining $h$ (Equation 1) as the percept-to-stimulus mapping makes its inverse function $h^{-1}$ easier to train, as shown in Figure 5, where the MSE losses are 0.2001 vs. 0.2756 at the step of 20000. Therefore, we report the results based on this setup in Table 2.
Conditional invertible neural network. The cINN architecture follows the design of the INN, with the additional conditioning to the subfunction at each layer [4, 2]. The percept or the MNIST target is concatenated as the conditioning during the training or the inference, respectively (Figure 3(b)). The conditional input $\mathbf{c}$, which is $\mathbf{y_{p}}$ during the training and $\tilde{\mathbf{x}}$ during the evaluation, is first transformed with a learnable linear layer to a size of $9\times 9$ before being passed to the subfunction. As introduced in Section 3.3, the cINN is trained in an unsupervised manner only with the NLL loss (Equation 7).
Table 2: Reconstruction quality evaluation from the predicted percepts to the original visual signals based on the optimized stimuli. Five approaches are compared with two resolutions (Res.) of the original MNIST input image of $9\times 9$ and $28\times 28$ pixels. Five evaluation metrics are reported, namely Mean Absolute Error (MAE), Mean Squared Error (MSE), Structural Similarity Index Measure (SSIM), Peak Signal-to-Noise Ratio (PSNR), and classification accuracy with a pre-trained classifier (ACC). The number of the parameters of each model is listed. The top results are marked in blue, while the second-best results are marked in red.
 Methods    Res. MAE $\downarrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ ACC $\uparrow$ #Param $\downarrow$
down 0.1682 0.0619 0.7248 30.4149 0.4272 1
linear 9 0.1657 0.0584 0.7246 30.6877 0.5608 6,561
NN $\times$ 0.1387 0.0436 0.7799 31.9620 0.5572 946,512
INN 9 0.1785 0.0599 0.7088 30.5656 0.4892 1,866,270
cINN 0.1373 0.0570 0.7651 30.8940 0.5480 1,936,710
down 0.3432 0.2040 0.3784 25.0973 0.2884 1
linear 28 0.3279 0.1999 0.4770 25.1453 0.5884 63,504
NN $\times$ 0.3340 0.1966 0.4901 25.2400 0.6192 4,270,900
INN 28 0.3231 0.1947 0.4937 25.2729 0.5764 18,063,390
cINN 0.3136 0.1845 0.4943 25.5161 0.5932 2,449,197
(a) Examples of random samples from $9\times 9$ stimuli to both $9\times 9$ and $28\times 28$ percepts using Axon Map Model [8] implemented in pulse2percept [7].
(b) Visualization of the reconstruction quality with five approaches. (I) denotes the original MNIST targets. (II) denotes the predicted percepts of $28\times 28$ pixels based on (III) optimized stimuli of $9\times 9$ electrodes.
Figure 7: Visualization during the training and the inference.
@@ 4.4 Likelihood and MSE
We illustrate whether the maximum likelihood training in the cINN approach leads to decrease the MSE loss, as it is reported that good likelihood does not always guarantee good samples in generative models [52]. The likelihood of the optimized stimulus depends on the likelihood of the latent $\mathbf{z}$ (Equation 3). Figure 6(a) shows the relationship between the log-likelihood of optimized stimulus (with $\mathbf{z}=\mathbf{0}$) and the MSE between the predicted percept and MNIST target (Figure 3(b)). For both random and MNIST samples, the correlation between the log-likelihood and the MSE can be observed. While setting $\mathbf{z}=\mathbf{0}$ should already give a good result, we further investigate if better $\mathbf{z}\neq\mathbf{0}$ can be found to yield higher likelihood and thus lower MSE. We select the worst 50 MNIST examples with the highest MSE and modify the latent $\mathbf{z}$ to maximize the total likelihood of optimized stimulus using gradient descent. Figure 6(b) shows that the MSE values are indeed decreased for most of the samples. Further closed-loop optimization w.r.t. $\mathbf{z}$ can be addressed in future work (e.g., [24, 37]).
@@ 4.5 Reconstruction Quality
The visualization of the reconstruction quality with a percept resolution of $28\times 28$ pixels are illustrated in Figure 7. We argue that negative stimulations (in red) can counteract the surrounding or elongated phosphenes, which is not applicable in downsampling. The predicted percepts (II) from the cINN show clearer turnings (e.g., the bottom left part in the image of the digit ’2’) and cleaner corners of the visual field. The advantages can be confirmed by quantitative assessments. Five evaluation metrics are reported in Table 2, namely Mean Absolute Error (MAE), Mean Squared Error (MSE), Structural Similarity Index Measure (SSIM) [58], Peak Signal-to-Noise Ratio (PSNR), and classification accuracy with a pre-trained classifier (ACC) [19, 62]. These metrics encompass pixel-level measurements (MAE, MSE), fidelity-oriented assessments (SSIM, PSNR), and semantic evaluations (ACC). The classifier is pre-trained on the MNIST dataset [32], achieving a validation accuracy of 0.9921 for downsampled $9\times 9$ pixel images and 0.9954 for the original $28\times 28$ pixel images. As shown in Table 2, cINNs demonstrate strong performance in both input resolutions, balancing pixel-level reconstruction with image-level semantic preservation.
4.6 Resolution Discrepancies
We experiment with two different visual signal resolutions ($9\times 9$ and $28\times 28$ pixels) using a fixed number of electrodes ($9\times 9$). Resolution discrepancies between the stimulus and the percept indicate that retinal prosthetic optimization is fundamentally a non-bijective inverse problem. With an increased percept resolution, cINN outperforms other approaches in various metrics. The number of the parameters of each model is listed in Table 2. The supervised models (linear model, feed-forward neural network and the MSE part in invertible neural networks) are susceptible to a higher resolution due to the fully-connected layers in the architectures, while this dense layer is only used to concatenate the target or percept to the subfunctions in the cINN approach. Consequently, the parameter count increases by only 1.26 times when using cINNs, compared to a 9.68-fold increase when using INNs with a corresponding 9.68-fold increase in the input size (Table 2). Thanks to their inherent invertibility and generativity, both INN-MMD and cINN show robust results when the input resolution is approximately 10 times larger than the electrode array of the retinal implant. Although deep neural networks have a large number of parameters, their implementation should not be problematic on modern portable hardware, potentially leading to substantial performance improvements of reconstruction quality.
@@ 5.1 Limitations
Our work relies on a pre-defined Axon Map Model with fixed $\rho$ and $\lambda$ for simulation. However, these parameters are largely patient- and implant-dependent [9] and not facile to determine in real-world scenarios (cf. [16, 18]). Additionally, the pre-computed random samples, as illustrated in Figure 6(a), are not sufficiently general, making the study of more complex datasets a challenge. Finally, we did not compare our quantitative results with those of other papers in their original forms, as they are based on different assumptions and setups, such as the computational model used in retinal prosthetic simulation and the number of electrodes in the retinal implant. Nevertheless, we contend that our experiments include a diverse range of approaches, demonstrating the proof-of-concept for leveraging INNs in retinal prosthetic optimization.
@@ 5.2 Outlook
We anticipate our future studies to advance in three directions. First, supervised training of INN [3] can be further explored using different invertible architecture designs and training strategies. Second, simulation experiments can be conducted on more complex datasets utilizing other computational models to optimize stimuli. Third, eye movements of human beings, such as fixations, saccades, and micro-saccades, can be incorporated into the simulation and optimization processes.